# Simple Network Mechanism Leads to Quasi-Real Brain Activation Patterns with Drosophila Connectome

### v1

@@ Abstract
Considering the high computational demands of most methods, using network communication models to simulate the brain is a more economical way. However, despite numerous brain network communication models, there is still insufficient evidence that they can effectively replicate the real activation patterns of the brain. Moreover, it remains unclear whether actual network structures are crucial in simulating intelligence. Addressing these issues, we propose a large scale network communication model based on simple rules and design criteria to assess the differences between network models and real situations. We conduct research on the biggest adult Drosophila connectome data set. Experimental results show significant activation in neurons that should respond to stimulus and slight activation in irrelevant ones, which we call quasi-real activation pattern. Besides, when we change the network structure, the quasi-activation patterns disappear. Interestingly, activation regions have shorter network distances to their input neurons, implying that the network structure (not spatial distance) is the core to form brain functionality. In addition, giving the input neurons a unilateral stimulus, we observe a bilateral response, which is consistent with reality. Then we find that both hemispheres have extremely similar statistical indicators. We also develop real-time 3D large spatial network visualization software to observe and document experimental phenomena, filling the software gap. This research reveals network models power: it can reach the quasi-activation pattern even with simple propagation rules. Besides, it provides evidence that network structure matters in brain activity pattern generation. Future research could fully simulate brain behavior through network models, paving the way for artificial intelligence by developing new propagation rules and optimizing link weights.
@@ 1 Introduction
The neural systems of intelligent beings maintain constant communication[1][2]. The intricate mechanism of this communication unfolds as follows: specific neurons, often input neurons, receive stimulus signals from the external environment or other parts of the body, transitioning from a resting to an activated state. Subsequently, they relay these stimulus signals to the next tier of neurons through synapses. This process gives rise to macroscopic emergence across the entire neural system. Such communication fosters the emergence of intelligence in all intelligent life forms, including humans. Consequently, comprehending and replicating this process has become an indispensable avenue for unraveling biological intelligence and realizing artificial intelligence[3].
@@
Neural systems are interconnected networks[4], with neurons engaging in constant communication via synapses within the network’s topology. The brain’s unique and complex network structure is pivotal to the emergence of intelligence[5]. Many scientists posit that network science is crucial for deciphering the mechanisms underlying the flexible regulation of neural communication[6][7][8]. Determining how to leverage network science methodologies stands as a paramount challenge in contemporary neuroscience. Within this discipline, neural networks are often represented as graphs, consisting of nodes (entities) and edges (relationships). From the broadcasting model[9] to the diffusion model[10], researchers strive to craft network models that mimic activation patterns observed in the brain, offering insights into the workings of neural networks.
However, despite artificial neural networks (ANNs)’ deviation from the brain’s actual structure, they get full achievements in achieving intelligence, which cast doubt on the necessity of real network structures in brain behavior simulation. Furthermore, some researchers have highlighted the absence of evidence linking network model simulations to genuine brain activity patterns[11], undermining the models’ status as a robust research tool in neuroscience. The effectiveness of network models in replicating the neuronal system’s activation dynamics remains uncertain. This uncertainty partly stems from many models’ failure to replicate the brain’s structure at the neuron and synapse level, compromising the models’ interpretability.
@@
Moreover, current brain network communication models often rely on prevalent metrics such as signaling cost and activation time to assess model quality[11]. Yet, these metrics fall short of verifying whether a network model accurately reflects the brain’s actual activation patterns, which is the fundamental goal of brain simulation. There is a need for evaluation criteria that can affirm whether network communication models genuinely mirror the brain’s communication processes. Additionally, observing the brain network communication process is important for researches. The absence of 3D visualization software capable of real-time monitoring of communication status and displaying networks with a vast number of nodes poses challenges for researchers, including neuroscientists, by impeding real-time observation, display, and analysis of experimental outcomes.
Over the past few decades, biologists have dedicated themselves to determining the precise structure of the connectome, ranging from C. elegans[12] to Drosophila larvae[13]. By understanding the connections within an organism’s nervous system, researchers can gain insights into how the nervous system operates. A recent study[14] has published the largest adult Drosophila connectome, providing data support for researches in network science.
For these reasons, we conducted experiments utilizing the recently released comprehensive data set of the largest adult Drosophila connectome, enhancing the interpretability of our findings. Our approach involves crafting a large scale network communication model based on simple mechanisms, designed to simulate the activation behavior observed in the connectome.
Specifically, We establish the network representation of the connectome. Then we design each neuron’s state and threshold functions. When a neuron’s state meets the threshold, it activates and subsequently transmits signals to other neurons. After a few iteration, the whole network emerges an activation pattern. Additionally, we propose an evaluation criteria based on the ratio of actually activated neurons to the expected number of activation. Experiments reveals that our model has significant activation in the type of neurons that should be affected by the stimulus and slight activation in irrelevant neurons which called a quasi-real activation pattern. To prove the vitalness of network structure, we calculate every node’s layer and shuffle part of the network structure in same layer and find the shuffled network can’t generate any quasi-real activation pattern.
To enhance our understanding, we developed a 3D network visualization software, utilizing HTML5’s Canvas and Three.js based on WebGL technology for large spatial network visualization effects. This tool enables real-time observation of each neuron’s communication process.
@@
To prove the network structure matters (rather than spatial structure) in activation pattern generation, we calculate the average network and spatial distance of input and activation neurons and results reveal that the average network distance between input neurons and their corresponding activation areas is shorter than unrelated areas, which is a strong evidence to prove network structure matters in the process of message transmission in brain.
@@
Furthermore, by subjecting input neurons to the same stimulus, we observe significant responses in neurons expected to be affected, with only slight responses in irrelevant neurons—characterized as a quasi-real activation pattern. Notably, when providing unilateral stimuli to input neurons, we observe a bilateral response in the Drosophila neuronal wiring diagram, aligning with reality. Then we calculate the statistics properties and find that the left and right hemispheres have extremely similar degree, clustering coefficient, and eigenvector centrality vector coefficient. The vectors’ Person coefficient are up to 0.9986, 0.9988, and 0.9991, respectively.
2 Related Works
Various models have been developed to mimic or understand aspects of brain function, ranging from the cellular level to the whole brain. In this section we review the models that aim to simulate the brain’s behavior. The review includes neural dynamic model that describes how action potentials in neurons are initiated and propagated. In addition, we introduce the large scale brain simulating methods. At last, we introduce the network communication models which refers to a computational or theoretical framework designed to understand and simulate how different parts of the brain communicate with each other.
2.1 Neuron Dynamic Model
In past years, neuroscientists have aimed to describing the behavior of neurons by dynamic model. One of the most famous works is the Hodgkin-Huxley (HH) model[15]. It includes a set of differential equations that describe the changes in membrane potential and the dynamics of ion channels. The key variables in the model are the membrane potential $V$, and the gating variables $m$, $n$, $h$ that represent the probability of ion channels being open.
The currents are given by:
$I_{Na}$ $=\bar{g}{Na}m^{3}h(V-E{Na})$ (1)
$I_{K}$ $=\bar{g}_{K}n^{4}(V-E_{K})$ (2)
$I_{L}$ $=\bar{g}_{L}(V-E_{L})$ (3)
$\frac{dm}{dt}$ $=\alpha_{m}(1-m)-\beta_{m}m$ (4)
$\frac{dn}{dt}$ $=\alpha_{n}(1-n)-\beta_{n}n$ (5)
$\frac{dh}{dt}$ $=\alpha_{h}(1-h)-\beta_{h}h$ (6)
Another famous theory is cable theory[16]. Cable theory is a mathematical model used to understand how electrical signals propagate along neurons, especially in their dendrites and axons. This theory applies the principles of electrical circuits to describe how the passive properties of the neuron influence signal transmission. It’s particularly useful in explaining the conduction of graded potentials, such as synaptic potentials and electrotonic (passive) spread of membrane potential changes.
$\frac{\partial^{2}V}{\partial x^{2}}=\frac{1}{r_{m}}\left(\frac{\partial V}{\partial t}+\frac{V}{r_{c}}\right)$ (7)
, where $V$ is the membrane potential, $x$ is the distance along the cable, $t$ is time, $r_{m}$ is the membrane resistance per unit length, $r_{c}$ is the cytoplasmic (axial) resistance per unit length. This equation describes how the voltage $V$ changes along the length of the neuron $x$ and over time $t$.
Cable theory has several important applications in neuroscience, including understanding signal decay, neural modeling and synaptic integration. Despite its usefulness, cable theory has limitations. First, real neurons have complex geometries, which are simplified in cable theory. This can lead to inaccuracies in predicting real neuronal behavior. The complex structure of neuron should not be neglected. Besides, cable theory primarily addresses passive electrical properties and does not account for active properties like action potentials, which are crucial for neural communication.
One of the simplest and most commonly used models to describe neuron dynamics in SNNs[19] is the Leaky Integrate-and-Fire (LIF) model[20]. The LIF model captures the essential characteristics of spike generation in neurons. The membrane potential $V(t)$ of a neuron at time $t$ is described by the following differential equation.
$\tau_{m}\frac{dV(t)}{dt}=-[V(t)-V_{rest}]+RI(t)$ (8)
Despite their advances, SNNs face significant challenges. One primary issue is that the SNN’s structure often draws inspiration from anatomically defined from functional brain networks rather than accurately replicating the intricate neural connections observed at the microscale. In terms of the connection method of artificial neurons, SNNs usually use a structure similar to that of ANNs. Consequently, these models resort to architectures akin to those found in artificial neural networks, hindering their ability to authentically emulate brain functions. Additionally, SNNs are marked by their low computational efficiency, with a single forward propagation frequently requiring several seconds to complete. These limitations underscore the difficulties in considering SNNs as the optimal approach for brain simulation in their current state.
As for the large scale brain simulation model aims to simulate the whole brain’s behavior, Eugene et al.[25] present the first large-scale brain simulation model based on experimental measures in several mammalian species. The model is founded upon the global thalamocortical anatomy derived from diffusion tensor imaging of the human brain. It incorporates multiple thalamic nuclei and a six-layered cortical microcircuitry, which is grounded in in vitro labeling and three-dimensional reconstruction of single neurons from the cat visual cortex. Additionally, the model simulates over 1,000,000 SNN neurons, calibrated to replicate known in vitro response patterns recorded in rats. It boasts nearly half a billion synapses, incorporating accurate receptor kinetics, short-term plasticity, and dendritic spike-timing-dependent synaptic plasticity (dendritic STDP). The model can demonstrate behavioral patterns that mimic normal brain activity. This model is a comprehensive attempt to simulate the whole neural system, which is a critical component of the brain involved in sensory processing and consciousness.
Eliasmith et al.[26] introduce a more comprehensive brain simulating model, which consists of 2,500,000 neurons and serves to close this divide by demonstrating a variety of behaviors. This model receives only sequences of visual images and produces all its responses through a simulated arm. Despite its simplifications of real structure, the model encapsulates numerous elements of neuroanatomy, neurophysiology, and behavioral psychology, as evidenced by its performance across eight distinct tasks.
As for the standardized platform, Zeng et al. [27] present BrainCog, a cognitive intelligence engine inspired by the brain. This platform, based on SNN, offers critical infrastructure for the creation of brain-inspired artificial intelligence and simulation of the brain. BrainCog encompasses a variety of biological neurons, encoding methods, learning principles, brain regions, and a combination of hardware and software as its core elements.
However, it is essential to note that each neuron’s dynamic model involves multiple differential equations, posing a significant challenge in simulating the complete behavior of the intact brain given the constraints of available computing power. Therefore, there is a compelling need for a cost-effective simulation approach to address this challenge.
2.2 Brain Network Communication Model
The network communication model is a conceptual framework that describes how data is transmitted from one node to another over a network. It breaks down the complex process of communication into simpler, more manageable layers or components, each with a specific role. The most widely recognized network communication models are the Open Systems Interconnection (OSI) model and the Internet Protocol Suite, commonly known as the TCP/IP model. Nowadays, scientists apply network communication models to neuron research, which are mostly used to understand how brain networks operate[30][31][32].
Early studies have indicated that, similar to social and electrical networks[33], brain networks also exhibit properties such as power-law distributions[34] and the rich-club phenomenon[35]. These results indicate that brain networks contain a wealth of network statistical characteristics. To reveal the mystery of the brain, researchers must take network characteristics into consideration.
Researchers designed many brain network models to describe the processes of message propagation in brain. One category of brain network communication models is the routing protocol model. The most famous model of this category is shortest path model[36], which assumes that signals between a pair of nodes travel along the shortest path between them. Let $d(v)$ represent the shortest path length from the starting node to node $v$, for each node $v$ in the network, initialize $d(v)=\infty$, then for the starting node $s$, set $d(s)=0$. Next, process each node in increasing order of $d(v)$, for each node, update the $d(u)$ value of its neighbors $u$. The update way can be represented as follows:
$d(u)=\min(d(u),d(v)+w(v,u))$ (9)
where $w(v,u)$ is the weight of edge $e_{uv}$.
However, the key disadvantage of the model is these methods require each neuron must have knowledge about the entire network structure[37], which doesn’t correspond to the reality: every neuron only has the message of its neighbor neurons.
@@ 2.2 Brain Network Communication Model
However, the main disadvantage of navigation model is it supposes that neurons know the distance between their neighbors and a desired target region. Recently their still has a lack of evidences that reveal how this knowledge would be given to neurons in brain networks.
$P_{ij}=\frac{1}{k_{i}}$ (10)
$P_{ij}=\frac{w_{ij}}{\sum_{l\in N(i)}w_{il}}$ (11)
Where $w_{ij}$ is the weight of node $i$ to node $j$,$N(i)$is the set of node $i$.
However, studies have shown that the efficiency of signal transmission through random walks is relatively low[41]. So it can’t describe the rapid communication process in the brain.
$a_{i}(t+1)=1\quad\text{if}\quad\frac{1}{k_{i}}\sum_{j\in N_{i}}a_{j}(t)\geq\theta_{i}$ (12)
Where $N_{i}$ is the set of neighbors of node $i$, $a_{i}(t)$ is the adoption state of node $i$ at time $t$. $\theta_{i}$ is the threshold of node $i$. $k_{i}$ is the degree of node $i$. The threshold model has been extensively applied in artificial neural networks.
However, how to set the propagation thresholds for different neurons, and how to design the dynamics model between neurons remains a challenge.
3 Data and Methods
3.1 Drosophila Connectome Data Set
We employ the recently released largest Drosophila brain connectome[14], offering a thorough mapping of the entire neural network within a Drosophila female brain. This data set encompasses over 120,000 neurons and 30,000,000 synapses, including detailed information on synaptic connections between neurons. Notably, it provides high-precision 3D coordinates, accurate to the nanometer level, for both cell bodies and synapses. Additionally, this data set offers extensive labeling for various types of neurons. For a more in-depth understanding, please refer to [14][43] for detailed information.
Table 1 shows the basic properties of the data set. Note that we combine multiple edges between two nodes to calculate network statistic properties. The average degree before combining multiple edges is 501.6 while average degree is 37.01 after combination. It means most of the pairs of neuron has many links.
Number of Nodes Number of Edges Average Degree
$131459$ $32970606$ $37.01$
$501.6$ $0.1527$ $0.0008817$
Apart from that, we draw the curves of degree distribution. From Fig. 1 We found that the distribution of log(degree) initially appears to be approximately linear, but at the end it exhibits a long-tail distribution, which demonstrate that the original distribution of degree follows a power-law distribution. This result corresponds to the earlier research that neural network also has the power-law property[34].
Figure 1: Degree Distribution (Log-Log Scale)
Also, we plot the correlation curves between degree and cell surface area, volume, and cell length (See in Fig. 2). The results indicate the degree distribution is logarithmically linearly correlated with these data. These characteristics of the brain can explain Weber-Fechner’s law[44]: for any sensory pattern, the perceived intensity is a logarithmic function of the physical intensity. The structural changes can only cause a logarithmic change in the average connection strength.
Figure 2: Correlation between degree distribution and neuron characteristics

### v3

# Network Structure Trumps Neuron Dynamics: Insights from Drosophila Connectome Simulations

@@ Abstract
Despite the success of artificial neural networks, the necessity of real network structures in simulating intelligence remains unclear. Utilizing the largest adult Drosophila connectome data set, we constructed a large-scale network communication model framework based on simple neuronal activation mechanisms to simulate the activation behavior observed in the connectome. The results demonstrate that even with simple propagation rules, models based on real neural network structures can generate activation patterns similar to those in the actual brain. Importantly, we found that using different neuronal dynamics models, all produced similar activation patterns. This consistency across different models emphasizes the crucial role of network topology in neural information processing, challenging views that rely solely on neuron count or complex individual neuron dynamics. Moreover, we test the influence of network reconnect rate and find that at least 1%’s reconnect rate will ruin the activation patterns appeared before. By comparing network distances and spatial distances, we found that network distance better explains the information propagation patterns between neurons, highlighting the importance of topological structure in neural information processing. To facilitate these studies, we developed real-time 3D large spatial network visualization software, bridging a crucial gap in existing tools. Our findings underscore the importance of network structure in neural activation and provide new insights into the fundamental principles governing brain functionality.
@@ 1 Introduction
The neural systems of intelligent beings maintain constant communication[1, 2]. The intricate mechanism of this communication unfolds as follows: specific neurons, often input neurons, receive stimulus signals from the external environment or other parts of the body, changing from a resting to an activated state. Subsequently, they relay these stimulus signals to the next tier of neurons through synapses. This process gives rise to macroscopic emergence across the entire neural system. Such communication fosters the emergence of intelligence in all intelligent life forms, including human beings. Consequently, comprehending and replicating this process has become an indispensable avenue for unraveling biological intelligence and realizing artificial intelligence[3].
@@
Neural systems are interconnected networks[4], with neurons engaging in constant communication via synapses within the network’s topology. Some scientists think the brain’s unique and complex network structure is pivotal to the emergence of intelligence[5]. Many scientists posit that network science is crucial for deciphering the mechanisms underlying the flexible regulation of neural communication[6, 7, 8]. Determining how to leverage network science methodologies stands as a paramount challenge in contemporary neuroscience.
However, whether this network structure determines the emergence of intelligence remains a question. In recent years, artificial neural networks (ANNs) have taken rapid development in enormous fields[9, 10, 11]. A notable recent achievement is large language model (LLM). However, the structure of any ANNs does not follow the actual brain network structure strictly, which poses a challenge to the viewpoint that network structure inherently contains intelligence. Simultaneously, many scientists in machine learning adhere to the scaling law: the more parameters the network has, the greater the degree of emergent intelligence[12]. This viewpoint suggests that intelligence is solely related to the number of parameters a model has, rather than the unique network structure between neurons. Under this perspective, the reason biological brains generate intelligence is due to the number of neurons, and the unique network structure is a coincidence of evolution.
To demonstrate the importance of brain network structure, network scientists are constantly exploring and experimenting. An approach is brain network communication model[13]. Scientists wants to use brain network communication model to understand and describe the interaction and communication between different parts of the brain. This model aids in studying the structural and functional aspects of neuronal networks and how they coordinate to accomplish complex cognitive processes[14].
The first challenge of this approach is obtaining a complete and accurate mapping of the brain’s structure. For example, current imaging technologies, such as MRI, fMRI, and electron microscopy, have limitations in resolution and speed. While MRI and fMRI provide non-invasive imaging, their spatial resolution is not sufficient to map individual neurons and synapses. On the other hand, techniques like electron microscopy, which can achieve higher resolution, are extremely slow and generate enormous amounts of data that are challenging to process[15]. Modeling from the micro scale is necessary because modeling at the level of brain regions or nerve fibers does not include the most basic structural information and lacks the most complete physical foundation.
@@
Moreover, current models often rely on prevalent metrics such as signaling cost and activation time to assess model quality[16]. Yet, these metrics fall short of verifying whether a network model accurately reflects the brain’s actual activation patterns, which is the fundamental goal of brain simulation. There is a need for evaluation criteria that can affirm whether network communication models genuinely mirror the brain’s communication processes.
Based on facts and challenges above, benefiting from a largest-to-date adult Drosophila connectome data[17], this study attempts to investigate how the type of structure of neurons could affect the activation patterns of a brain. Our approach involves crafting a large scale network communication model framework based on simple mechanisms, designed to simulate the activation behavior observed in the connectome. Specifically, we establish the network representation of the connectome. Then we design a large scale threshold propagation model of nerve cells: when a neuron received enough stimuli, it activates and subsequently transmits signals to other neurons. After a few iterations, the whole network emerges an activation pattern. We use threshold model and sigmoid model as the basic neuron model. We also design an improved LIF model as the basic neuron model for comparison. Additionally, we propose an evaluation criteria based on the ratio of actually activated neurons to the expected number of activation. Experiments reveals that, from threshold , sigmoid, and LIF neuron model, no matter activation mechanism is, our model has significant activation in the type of neurons that should be affected by the stimulus and slight activation in irrelevant neurons. Moreover, to supply more evidence of the vitalness of network structure, we disturb the network by the way of degree-preserving rewiring and and find the disturbed network can’t generate any correct activation pattern.
In addition, we test the sensitivity of disturb degree. For details, we record the activation rate of areas under different reconnect rates and find that 1%’s random change of network structure will cause a complete change in function.
@@
To prove the network structure, rather than the spatial structure, matters in activation pattern generation, we calculate the average network and spatial distance of input and activation neurons and results reveal that the average network distance between input neurons and their corresponding activation areas is shorter than unrelated areas, which is a strong evidence of network matters in message propagation in brain.
@@
Furthermore, when providing unilateral stimuli to input neurons, we observe a bilateral response in the Drosophila connectome, aligning with reality. Then we calculate the statistics properties and find that the left and right hemispheres have extremely similar degree, clustering coefficient, and eigenvector centrality vector coefficient. The vectors’ Pearson coefficient are up to 0.9986, 0.9988, and 0.9991, respectively.
We also developed a 3D network visualization software see in appendix F, utilizing HTML5’s Canvas and Three.js based on WebGL technology for large spatial network visualization effects. This tool enables real-time observation of each neuron’s communication process.
2 Results
First, we report the basic statistics of Drosophila connectome. After that, to test the performance of our model in terms of realism on the data set and determine whether different artificial mechanisms can achieve similar activation patterns when the network structure changes, we give different stimuli and record the activation pattern of different neurons. In addition, we give a unilateral stimulus (stimulus one hemisphere) and record the bilateral response of neurons. Besides, to add more neurons’ real dynamic details, we also add an improved LIF model as the basic neuron dynamic model and record the activation patterns under stimuli. Finally, to prove the network structure matters in brain network communication, we calculate the average physical and network distance between different areas.
2.1 Experimental Settings
The basic methodology of the experiment is described as follows. In this study, a continuous stimulus signal (binary 0-1 signal) is applied to a subset of input neurons (such as visual or olfactory neurons) in the brain. Subsequently, we compute the signal’s propagation process within the brain and record the responses of both intermediate and output neurons. The evaluation criterion is determined by calculating the ratio of neurons that ought to be activated to those that are actually activated in a specific area (see as Equ.26 in appendix B). Given that the brain functions as a type of shallow neural network[53], we document the outcomes at an iteration step size of 5. Moreover, as the input remains constant and no randomization parameters have been introduced, the outcomes of each experiment are consistent, eliminating the need for repetition. For messages in related neurons, please read appendix C.
In linear threshold model, we set $\sigma_{0}=0.8$. In the simulation of the LIF model, numerous hyperparameters need to be used. All the hyperparameters used are listed in the following table. Biological hyperparameters are set according to real biologic experiment test[54]. Current proportion const is the optimal parameter obtained under equidistant search.
Table 1: List of Hyperparameters
Hyperparameter Description Value
$\tau$ Membrane time constant 15.0 ms
$V_{\text{rest}}$ Resting membrane potential -65 mV
$V_{\text{th}}$ Spike threshold -45 mV
@@ 2.1 Experimental Settings
$dt$ time step 0.55 ms
$\alpha$ Current proportion const 0.03/0.045 (visual/olfactory)
2.2 Visualization and Statistic Properties of Drosophila Connectome
In this section we first present the visualization of Drosophila Connectome (Fig.1(a)). Other visualization cases and methods of visualization are presented in appendix F. Visually, the shape of the brain appears to be physically similar, but whether the network structures are similar remains unknown. So we further investigate the network statistics of two hemispheres. We calculated the degree, clustering coefficient, and eigenvector centrality distribution for every node in each hemisphere (Fig.1(d), (e) and (f)). The Pearson correlation coefficient of degree, clustering coefficient, and eigenvector centrality vector in each hemisphere is 0.9986, 0.9988, and 0.9991, respectively. According to Fig.1(d), (e) and (f), these statistical indicators for both hemispheres are extremely similar, indicating that the brain exhibits strong symmetry in network structure.
From Fig.1(c) we can see that the degree, i.e., the number of connections to the nodes in the network, exhibits a long-tail distribution, implying that while a large number of nodes possess few connections, a small number of nodes contains a large number of connections[37]. It implies the real neural network have some same statistic properties as social, electrical and transportation networks. Also, we plot the correlation curves between degree and cell surface area, volume, and cell length (Fig.1). The results indicate that the degree distribution is logarithmically linearly correlated with these characteristics. This logarithmic relationship reveals a fundamental principle in the organization of neural networks in the brain.
To elaborate, as the surface area, volume, or length of a neuron increases, the number of synapses (represented by the degree) it forms grows logarithmically rather than linearly. This non-linear scaling is crucial for maintaining efficient information processing in the brain while preventing an explosion in connectivity that would be computationally and energetically costly.
These characteristics of the brain can explain Weber-Fechner’s law[47]: for any sensory pattern, the perceived intensity is a logarithmic function of the physical intensity. Specifically, the size, length, and area changes can only cause a logarithmic change in the number of synapses the neuron has. This means that as a stimulus becomes stronger, the increase in neural activity or connectivity is proportionally smaller, mirroring the logarithmic perception of intensity described by Weber-Fechner’s law.
(a) Visualization of Drosophila Connectome
(b) Correlation between degree distribution and neuron characteristics
Figure 1: Visualization and statistic properties of Drosophila connectome data set